Consider the following pairs of points for the variables X nd Y. Which one of the following represents the number of degrees of freedom that would be used in a test of hypothesis that the slope of the regression line between Y and X is different from zero?

x 0112345667 y66 745 23 011 From the data one obtains: SSG = 54.5, SS, = 54.5, SS, = -50.5, s, =.98148
Answer #1

n : number of pairs of data =10

number of degrees of freedom that would be used in a test of hypothesis that the slope of the regression line between Y and X is different from zero = n-2

Therefore,

Number of degrees of freedom = n-2 =10-2=8

Number of degrees of freedom = 8